OpenStudy (anonymous):

S6 for 2+1+1/2+1/4. I cant get it into sum form. Can you help me

OpenStudy (cwrw238):

this is a geometric series with common ratio 1/2 the sum of n terms Sn = a1 * (1 - r^n) ------- ( 1 - r) where a1 = first term and r = common ratio to get S6 plug in n=6, r=1/2 and a1 = 2

OpenStudy (cwrw238):

oh - they've gone!